Hi I am trying to figure out why I keep coming up with a (-2) for x. My book says the answer is a positive 2. Can someone please help me. Thanks.
p.s. I used the elimination method
3x=5y+11
2x=6y+2 Answer by josmiceli(19441) 
 3x=5y+11
2x=6y+2
x = 3y + 1
3x = 9y + 3
subtract from the 1st
0 = 4y -8
4y = 8
y = 2
2x = 6y + 2
2x = 6*2 + 2
x = 14/2
x = 7
check in the 1st
3x = 5*2 + 11
3x = 21
x = 7 checks